OpenStudy (anonymous):

INTRO TO PYTHON MIT 6.189 I'm looking at Exercise 2.3 of Homework 2 Handout PDF. I don't understand what I'm suppose to do for part 2. 2."...compute the following values using multadd and print out the results:" I have my multadd function but I don't get how the multadd function can be used to solve the equation listed

OpenStudy (e.mccormick):

You are looking at this: \(\sin\left( \dfrac{\pi}{4} \right) + \dfrac{\cos\left(\tfrac{\pi}{4} \right)}{2}\) \(\left\lceil \frac{276}{19} \right\rceil +2\log_7(12)\) and do not see: \(a\times b+c\) Did you ever think about the fact that due to order of operations: \(a\times b+c=c+a\times b\)